# MILD: Multi-Layer Diffusion Strategy for Complex and Precise Multi-IP Aware Human Erasing

### v1

@@ Abstract
Recent years have witnessed the remarkable success of diffusion models, especially in various image-customized tasks. Previous works have achieved notable advances on simple human-oriented erasing task by leveraging explicit mask guidance and semantic-aware inpainting paradigm. Despite this progress, existing methods still struggle with significant human removal challenges raised by “Multi-IP Interactions” under more complex realistic scenarios such as human-human occlusions, human-object entanglements and human-background interferences. These approaches are typically limited by: 1) Dataset deficiency. They basically lack large-scale multi-IP datasets, especially those covering complex scenarios such as dense occlusions, camouflaged or distracted backgrounds and diverse IP interactions. 2) Lack of spatial decoupling. Most methods lack effective spatial decoupling strategies to disentangle different foreground instances from the background, which limits their ability to achieve targeted human removal and clear background inpainting. These limitations result in performance degradation in precisely erasing overlapping individuals and restoring occluded regions in real-world scenes. In this work, we first introduce a high-quality human erasing dataset, MILD dataset, capturing diverse pose variations, occlusions, and complex backgrounds in multi-IP interactive scenarios. Building on this foundation, we propose Multi-Layer Diffusion (MILD), a novel multi-layer diffusion strategy for complex and precise multi-IP aware human erasing. Specifically, MILD first decomposes the traditional generation process into semantically separated denoising pathways, enabling independent reconstruction of each foreground instance and the background. To enhance human-centric understanding, we introduce Human Morphology Guidance, a plug-and-play guidance module that integrates pose, parsing, and spatial relations into the generation process, improving the morphological awareness and promoting more effective restoration. We further present Spatially-Modulated Attention, an adaptive mechanism that leverages spatial mask priors to effectively modulate attention across semantic regions, resulting in fewer boundary artifacts and mitigated semantic leakage. Extensive experiments demonstrate that MILD consistently outperforms state-of-the-art methods on challenging human erasing benchmarks. Our project page is available at: https://mild-multi-layer-diffusion.github.io/.
{strip}
Figure 1: The figure exhibits MILD (our method) can robustly handle three critical challenges in human erasing: Human–Human Occlusion, Human–Object Entanglement, and Human–Background interferences and can achieve clean and artifact-free removal results across diverse scenarios. Note each case shows the input and corresponding output pair.
@@ 1 Introduction
Object removal aims to eliminate user-specified regions and synthesize visually coherent, context-consistent backgrounds Jampani et al. (2021); Xu et al. (2023). This capability supports a wide range of applications, including photo editing, privacy preservation, and interactive content creation Yu et al. (2018); Liu et al. (2018). However, achieving both thorough removal and faithful background restoration remains a challenging task, especially in scenarios involving multiple interacting instances and dense occlusions.
@@
Early non-parametric methods Criminisi et al. (2004); Barnes et al. (2009) repainted erased regions by copying patches from visible areas, but often produce repetitive textures and fail in complex scenes. Afterwards, GAN-based models Pathak et al. (2016); Iizuka et al. (2017); Yu et al. (2018); Nazeri et al. (2019) introduce adversarial learning to enhance visual realism, yet suffer from global inconsistency and instability under large masks. Recently, diffusion-based models Song et al. (2021); Avrahami et al. (2023); Yang et al. (2023a) have achieved competitive performance by leveraging large-scale pretrained generative priors Rombach et al. (2022b); Ekin et al. (2024). These models Zhuang et al. (2024); Yildirim et al. (2023) excel in generating diverse content, but remain limited in eliminating objects in complex multi-instance scenarios. For example, in Figure 4, standard LDM always fails to remove the target cleanly and instead hallucinates a new human-like figure at the masked region. Notably, the generated content retains dominant visual hints, such as “human”, implying that semantic features from the removed subject have leaked into the generation process.
@@
These limitations highlight a fundamental challenge: most existing methods treat object removal as an unified generation task, without explicitly decoupling the generation pathways for different foreground instances. This framework often leads to semantic interference and content entanglement, especially in multi-IP or object-dense scenes, where features from one instance may bleed into others and degrade both the quality and controllability of the output.
To address these challenges, we introduce Multi-Layer Diffusion (MILD), a diffusion-based strategy that reformulates human erasing as a disentangled, layer-wise, instance-aware generation process. As illustrated in Figure 1, MILD achieves high-fidelity removal in three representative challenges: human-human occlusions, human-object entanglements, and human-background interferences, which represent typical scenarios where existing methods tend to perform suboptimally. Our contributions are threefold:
@@
  • MILD dataset. A high-quality human erasing dataset is curated capturing diverse scenes with challenging poses, occlusions, and human-instance interactions. This dataset provides paired images with precise removal masks and realistic background, enabling rigorous evaluation of human erasing models in real-world scenarios.
  • Multi-Layer Diffusion Strategy (MILD). We propose a MILD backbone that decomposes generation process into multiple semantically separated LoRA-based generation pathways, producing composable outputs corresponding to masked instances and the background. To further enhance performance, we introduce Human Morphology Guidance for instance-level understanding and Spatially-Modulated Attention for spatially conditioned attention control, enabling precise removal with minimal semantic leakage in complex scenes.
  • Comprehensive Experiments. Extensive experiments on both real-world and synthetic datasets evaluate the effectiveness of our MILD strategy using diverse quantitative metrics and human evaluations. The ablation studies further demonstrate the individual contributions of Human Morphology Guidance and Spatially-Modulated Attention. Collectively, the results confirm that MILD achieves state-of-the-art performance in removal accuracy and background fidelity, outperforming existing human erasing baselines.
Figure 2: (a) Overview of MILD. Given an input image and a set of target masks, the proposed Multi-Layer Diffusion (MILD) strategy performs human erasing by generating disentangled foreground layers and a background layer. A shared UNet with Layered LoRA enables efficient denoising across all branches, producing composable outputs. Spatially-Modulated Attention (SMA) injects adaptive biases to suppress semantic leakage across masked regions. Cross-attention is conditioned on the text prompt and Human Morphology Guidance (HMG) to guide instance-aware generation. (b) HMG extracts fine-grained human priors from pose and parsing maps using two dedicated encoders. These features, combined with a spatial mask prior, form a unified latent representation that enriches the denoising process with human-centric and spatial cues, improving fidelity and identity consistency.
@@ 2 Related Work
Image Inpainting
Image inpainting aims to restore missing or occluded regions in a visually coherent and semantically consistent manner in various scenarios Sargsyan et al. (2023); Zhou et al. (2020); Wang et al. (2024); Yoon and Cho (2024). Traditional methods relied on patch propagation Hays and Efros (2007); Ding et al. (2019), transferring low-level textures from visible regions but failing in semantically complex scenes. With the rise of deep learning, the encoder-decoder Zhou et al. (2023); Cao et al. (2023); Altinel et al. (2018); Li et al. (2020a); Yi et al. (2020a) and GAN-based architectures Iizuka et al. (2017); Yu et al. (2018); Zeng et al. (2019) have significantly improved structural plausibility and perceptual quality, further enhanced by edge guidance Nazeri et al. (2019) and attention mechanisms Yu et al. (2018). Diffusion-based methods Lugmayr et al. (2022); Saharia et al. (2022) have recently achieved strong generation performance by operating in latent spaces guided by powerful pretrained priors, but they often lack precise spatial control or instance-level disentanglement. Methods like SDXL Inpainting Podell et al. (2023) introduce mask-based conditioning for edit localization, but treating all masked regions as single entity prevents instance-specific control. Meanwhile, recent efforts incorporate large language models (LLMs) to enable instruction-based editing Brooks et al. (2023); Huang et al. (2024); Li et al. (2024b); Sheynin et al. (2024); Li et al. (2024c), yet they struggle to localize individual instances in complex scenes.
@@
Object removal, as a task derived from image inpainting, aims to eliminate user-specified content while preserving scene fidelity. Early approaches Suvorov et al. (2022) achieved reasonable results by adapting general-purpose generative models with task-specific Jain et al. (2023); Yi et al. (2020b) guidance, but they struggled to capture complex textures and scene structures. Recent works Hu et al. (2022); Kingma and Welling (2014); Ronneberger et al. (2015); Rombach et al. (2022a); Ho et al. (2020) have explored a range of strategies to improve controllability and fidelity Winter et al. (2024). Training-based Yang et al. (2023b); Jiang et al. (2025); Liu et al. (2025); Li et al. (2025) approaches fine-tune diffusion models using supervision such as masks, prompts, or parameter-efficient modules. Training-free Sun et al. (2025); Han et al. (2024); Jia et al. (2024); Li et al. (2024a) approaches, in contrast, manipulate internal representations (e.g., attention maps or embeddings) during inference to enable flexible editing. In parallel, different forms of user interaction have been introduced—ranging from text-prompted editing Bar-Tal et al. (2022); Zhao et al. (2021); Geng et al. (2024) to mask-guided Li et al. (2022); Köhler et al. (2014); Xie et al. (2023a, b) generation—to support more intuitive control. Despite these advances, we found that most existing methods lack instance-level distinction, generating a single prediction conditioned on all masked targets or prompts. This entangled treatment limits the compositional flexibility and often leads to interference between each instance and its surrounding region.
@@ 3 Method
Existing inpainting methods often fail to entirely remove a instance, leaving residual artifacts or introducing semantic leakage in complex scenes. To address these limitations, we introduce Multi-Layer Diffusion (MILD) (as shown in Figure 2), a diffusion-based strategy that reformulates human erasing as a layered generation process. MILD features a structured extension of low-rank adaptation as its backbone (Section 3.1) to produce disentangled per-instance and background layers. On top of this, Human Morphology Guidance (Section 3.2) injects pose, parsing, and spatial priors to enhance human-aware understanding, while Spatially-Modulated Attention (Section 3.3) dynamically suppresses semantic leakage across spatially separated regions with mask priors. Together, these components enable precise, artifact-free erasing in scenes with occlusions and entanglements.
3.1 MILD Backbone: Layered LoRA Diffusion
Traditional inpainting methods usually struggle with multi-IP scenes, as their unified generation process tends to produce residual artifacts and semantic confusion due to the lack of instance-level disentanglement. To address this limitation, we introduce a novel MILD backbone, a multi-branch adaptation of Low-Rank Adaptation (LoRA) that enables disentangled reconstruction of each instance while ensuring coherent scene composition.
Given an input image and $N$ target masks $\{M_{k}\}_{k=1}^{N}$, MILD Backbone produces $N$ foreground layers and a clean background layer. Each foreground branch is conditioned on its corresponding mask and reconstructs the associated subject, while the background branch synthesizes the unobstructed environment with all instances removed. To avoid redundant parameters and enable scalability, we adopt a shared UNet backbone with LoRA adapters. A single set of low-rank parameters is used across all foreground branches, while the background branch maintains an independent set. Formally, each LoRA modification is defined as:
$\Delta W_{p}^{(b)}$ $=\alpha_{p}^{(b)}\,B_{p}^{(b)}\,A_{p}^{(b)},$ $p\in\{Q,K,V\},b\in\{\mathrm{fg},\mathrm{bg}\},$ (1)
where $A_{p}^{(b)}\in\mathbb{R}^{r\times d}$, $B_{p}^{(b)}\in\mathbb{R}^{d\times r}$, and $\alpha_{p}^{(b)}$ is a learnable scaling factor that supports concurrent training and inference of all branches, and facilitates interaction between instance-level predictions and global background restoration.
During training, each foreground branch learns to reconstruct its assigned region by utilizing the loss:
$\mathcal{L}_{k}=\left\|M_{k}\odot\left(\epsilon-\epsilon_{k}\right)\right\|_{2}^{2},$ (2)
where $\epsilon\sim\mathcal{N}(0,1)$ is standard Gaussian noise and $\epsilon_{k}$ denotes the prediction from the $k$-th branch. The background branch predicts noise over the entire image using the loss:
$\mathcal{L}_{\text{bg}}=\left\|\epsilon-\epsilon_{\text{bg}}\right\|_{2}^{2}.$ (3)
The total loss is formulated by combining the foreground and background losses as follows:
$\mathcal{L}_{\text{total}}$ $=\lambda_{t}\sum_{k=1}^{N}\mathcal{L}_{k}+\mathcal{L}_{\text{bg}},$ (4)

### v2

# Multi-Layer Diffusion Strategy for Multi-IP Interaction-Aware Human Erasing

@@ Abstract
Recent years have witnessed the success of diffusion models in image customization tasks. However, existing mask-guided human erasing methods still struggle in complex scenarios such as human–human occlusion, human–object entanglement, and human–background interference, mainly due to the lack of large-scale multi-instance datasets and effective spatial decoupling to separate foreground from background. To bridge these gaps, we curate the MILD dataset capturing diverse poses, occlusions, and complex multi-instance interactions. We then define the Cross-Domain Attention Gap (CAG), an attention-gap metric to quantify semantic leakage. On top of these, we propose Multi-Layer Diffusion (MILD), which decomposes the generation process into independent denoising pathways, enabling separate reconstruction of each foreground instance and the background. To enhance human-centric understanding, we introduce Human Morphology Guidance, a plug-and-play module that incorporates pose, parsing, and spatial relationships into the diffusion process to improve structural awareness and restoration quality. Additionally, we present Spatially-Modulated Attention, an adaptive mechanism that leverages spatial mask priors to modulate attention across semantic regions, further widening the CAG to effectively minimize boundary artifacts and mitigate semantic leakage. Experiments show that MILD significantly outperforms existing methods. Datasets and code are publicly available at: project page.
Figure 1: The figure exhibits MILD (our method) can robustly handle three critical challenges in human erasing: Human–Human Occlusion, Human–Object Entanglement, and Human–Background interferences and can achieve clean and artifact-free removal results across diverse scenarios.
@@ 1 Introduction
Object removal aims to eliminate user-specified regions and synthesize visually coherent, context-consistent backgrounds (jampani2021slide; xu2023spectralhint). This capability supports a wide range of applications, including photo editing, privacy preservation, and interactive content creation (yu2018generative; liu2018image). However, achieving both thorough removal and faithful background restoration remains a challenging task, especially in scenarios involving multiple interacting instances and dense occlusions.
@@
Early non-parametric methods (criminisi2004region; barnes2009patchmatch) repainted erased regions by copying patches from visible areas, but often produce repetitive textures and fail in complex scenes. Afterwards, GAN-based models (pathak2016context; nazeri2019edgeconnect) introduce adversarial learning to enhance visual realism, yet suffer from global inconsistency and instability under large masks. Recently, diffusion-based models (song2021score; avrahami2023blended; yang2023diffedit) have achieved competitive performance by leveraging large-scale pretrained generative priors (rombach2022ldm; clipaway2023). These models (powerpaint2024; instinpaint2023) excel in generating diverse content, but remain limited in eliminating objects in complex human-centric multi-instance scenarios.
Figure 2: Peak signal of semantic leakage.
For example, in Figure 2, standard LDM always fails to remove the target cleanly and instead hallucinates a new human-like figure at the masked region. Notably, the generated content retains dominant visual hints, such as “human”, implying that semantic features from the removed subject have leaked into the generation process.
@@
These limitations reveal a fundamental challenge: most existing approaches treat object removal as a unified generation task without explicitly disentangling the generation paths of different foreground instances. This framework often leads to semantic interference and content entanglement, particularly in multi-interactive-person (Multi-IP) or object-dense scenes, where features from one instance may permeate into others, degrading output quality and controllability.
To address this challenge, we propose MILD (Multi-Layer Diffusion) strategy, a theoretically grounded framework for human-centric object removal. We first theoretically analyze the source of semantic leakage by introducing the Cross-Domain Attention Gap (CAG) to quantify attention flow discrepancies between foreground and background regions. Theoretical analysis demonstrates that enlarging CAG significantly reduces semantic leakage, inspiring our core design principle: maximizing the attention gap through architectural separation.
Guided by this theory, MILD reformulates human erasing as a layered diffusion process, establishing independent diffusion paths for each target instance and the background to achieve instance-level disentanglement. Specifically, MILD employs a shared UNet backbone with domain-specific LoRA adapters, ensuring scalability while maintaining attention domain separation between foreground and background. To enhance instance awareness in complex scenarios, we introduce Human Morphology Guidance (HMG), which injects human pose and parsing priors to preserve structural integrity under occlusion. Furthermore, the Spatially-Modulated Attention (SMA) module applies mask-conditioned constraints to attention logits, further expanding CAG and achieving exponential leakage suppression. This design not only enables more thorough instance removal but also produces flexibly recomposable layered outputs, supporting diverse scene editing operations.
@@
To advance research in this field, we release a high-quality human-erasing dataset with diverse scenes, challenging poses, occlusions, and interactions. It provides paired images with precise masks and realistic backgrounds for rigorous real-world evaluation of human-removal models.
Experimental results demonstrate that our MILD significantly outperforms existing methods on complex human removal tasks with superior visual realism, semantic consistency, and perceptual quality, confirming its state-of-the-art performance.
@@ 2 Related Work
We introduce the recent works concentrated in image inpainting (see Appendix  C), object Removal (see Sec.  2) and Layered Generation (see Appendix  C).
@@
Object removal, as a task derived from image inpainting, aims to eliminate user-specified content while preserving scene fidelity. Early approaches (suvorov2022lama) achieved reasonable results by adapting general-purpose generative models with task-specific (jain2023structuretexture; yi2020cra) guidance, but they struggled to capture complex textures and scene structures. Recent works (hu2021lora; kingma2014auto; ronneberger2015unet; rombach2022high; ho2020denoising) have explored a range of strategies to improve controllability and fidelity (winter2024objectdrop). Training-based (yang2023magicremover; smarteraser2025; erasediffusion2025; rorem2025) approaches fine-tune diffusion models using supervision such as masks, prompts, or parameter-efficient modules. Training-free (sun2025attentive; han2024proxedit; jia2024designedit; li2024source) approaches, in contrast, manipulate internal representations (e.g., attention maps or embeddings) during inference to enable flexible editing. In parallel, different forms of user interaction have been introduced—ranging from text-prompted editing (bartal2022text2live; zhao2021comodgan; geng2024instructdiffusion) to mask-guided (li2022mat; kohler2014maskspecific; xie2023smartbrush; xie2023dreaminpainter) generation—to support more intuitive control. Despite these advances, we found that most existing methods lack instance-level distinction, generating a single prediction conditioned on all masked targets or prompts. This entangled treatment limits the compositional flexibility and often leads to interference between each instance and its surrounding region, especially when it comes to complex human-oriented interaction scenes.
Figure 3: Overview of MILD. Given an input image and a set of target masks, the proposed Multi-Layer Diffusion (MILD) strategy performs human erasing by generating disentangled foreground layers and a background layer. A shared UNet with Layered LoRA enables efficient denoising across all branches, producing composable outputs. Spatially-Modulated Attention (SMA) injects adaptive biases to suppress semantic leakage across masked regions. Cross-attention is conditioned on the text prompt and Human Morphology Guidance (HMG) to guide instance-aware generation.
@@ 3 Method
Existing inpainting methods often fail to entirely remove an instance, leaving residual artifacts or introducing semantic leakage in complex scenes. To address these phenomenon, we define Cross-Domain Attention Gap (CAG) to quantify the semantic leakage. On the top of our theoritical analysis, we introduce Multi-Layer Diffusion (MILD) (Fig. 3), a diffusion-based strategy that reformulates human erasing as a layered diffusion process. MILD ultilizes low-rank adaptation modules as its backbone (Sec. 3.2) to produce disentangled per-instance and background layers. On top of this, Human Morphology Guidance (Sec. 3.3) injects pose, parsing, and spatial priors to enhance instance-aware understanding, while Spatially-Modulated Attention (Sec. 3.4) suppresses semantic leakage across spatially separated regions using mask priors. Together, these components enable precise, artifact-reduced erasing in scenes with occlusions and entanglements.
3.1 Theoretical Motivation: Disentanglement via Attention Gap
Traditional inpainting methods often produce residual artifacts and semantic confusion in multi-instance scenes, which attributing to the lack of the fundamental decoupling between foreground and background in unified generation processes. To address this issue, we first establish a theoretical framework for quantifying semantic leakage.
We introduce the Cross-Domain Attention Gap (CAG) as a measure of attention flow discrepancy between foreground and background regions. Our analysis reveals that semantic leakage correlates directly with the magnitude of this gap:
$\gamma:=\inf_{i\in B}\Biggl[\log\sum_{j\in B}e^{A_{ij}}-\log\sum_{j\in F}e^{A_{ij}}\Biggr],$ (1)
where $A_{ij}$ represents attention scores, and $B$ and $F$ denote background and foreground regions.
Theorem 1 (proof in Appendix F.3) establishes that a larger $\gamma$ leads to markedly reduced semantic leakage. This theoretical insight motivates our core design principle: maximizing the attention gap through architectural separation.
3.2 MILD Backbone: Multi-Layer Disentanglement
Building on the theoretical foundation above, we propose Multi-Layer Diffusion (MILD)—a diffusion-based strategy that reformulates human erasing as a layered diffusion process. Rather than relying on a unified single-layer pipeline that inherently couples instances, MILD implements the separation principle derived from our CAG analysis.
Given an input image and $N$ target masks $\{M_{k}\}_{k=1}^{N}$, MILD produces $N$ foreground layers and a clean background layer. Each foreground branch reconstructs its corresponding subject, while the background branch synthesizes the unobstructed environment. This architectural separation directly implements the attention gap maximization suggested by our theoretical analysis.
To ensure scalability while maintaining separation, we employ a shared UNet with domain-specific LoRA adapters. The parameterization follows our theoretical conditioning requirements:
$\Delta W_{p}^{(b)}=\alpha_{p}^{(b)}B_{p}^{(b)}A_{p}^{(b)},\quad p\in\{Q,K,V\},\quad b\in\{\text{fg},\text{bg}\},$
where foreground and background branches maintain independent adapter sets $(A_{p}^{(b)},B_{p}^{(b)})$, enforcing the attention domain separation identified as crucial for reducing semantic leakage.
The corollaries in Appendix F.4 and  F.5 shows that the multi-layer parameterization is better conditioned to train while being no worse in fit than a comparable single-layer model.
During inference, MILD produces composable outputs $\{\text{Layer}_{1},\ldots,\text{Layer}_{N},\text{Background}\}$ that can be flexibly recombined:
where $\mathcal{R}$ denotes instances to remove. This compositional approach directly embodies the decoupling benefits predicted by our CAG theory.
3.3 Human Morphology Guidance
Although the Layered LoRA diffusion backbone establishes the architectural foundation for instance disentanglement, complex scenarios involving spatial overlap and occlusion require enhanced instance-awareness to maintain structural fidelity. To handle such cases, we introduce Human Morphology Guidance (HMG) to reinforce instance identity under challenging conditions by injecting human-centric priors into the denoising process. HMG operates through two complementary guidance mechanisms: pose and parsing guidance and spatial context modeling. We use the same off-the-shelf estimators (cao2017realtime; li2020self) at training and inference pipeline and these cues are optional. For pose and parsing guidance, we extract pose keypoints and semantic parsing maps, which are encoded through shared convolutional stacks and linearly projected into compact morphological embeddings. These features are concatenated with the target instance’s mask embedding, providing explicit structural priors that help maintain body coherence under occlusion. To further enhance spatial reasoning, we add a lightweight mask-based context that explicitly models neighboring instances. For each instance $k$, we construct $M^{(k)}=\mathrm{clamp}\!\Bigl(\sum_{j\neq k}M_{j},\,0,1\Bigr),$ which aggregates all other instances into a single binary context. This explicit spatial modeling enhances localization accuracy and reduces interference between adjacent instances. As detailed in the algorithmic appendix (Appendix D), HMG seamlessly integrates into the MILD backbone by conditioning both foreground and background branches, with morphology features optionally incorporated during training and inference. The synergy between HMG’s explicit morphological guidance and our theoretical CAG framework yields significant benefits: explicit structural priors reduce the need for cross-instance attention, thereby amplifying the attention gap $\gamma$ identified in Section 3.1.
Figure 4: (a) HMG extracts fine-grained human priors from pose and parsing maps using two dedicated encoders. These features, combined with a spatial mask prior, form a unified latent representation that enriches the denoising process with human-centric and spatial cues, improving fidelity and identity consistency. (b) Illustration of the proposed Spatially-Modulated Attention (SMA) mechanism. For each query-key pair $(i,j)$, an spatial bias $\alpha_{st}$ is applied to the vanilla attention score $A_{ij}$, where $s=m_{i}$ and $t=m_{j}$ indicate their foreground/background status.
3.4 Spatially-Modulated Attention
Even though MILD backbone and HMG works to meet our decoupling design, we still observe residual semantic leakage due to limitations in the CAG under vanilla attention mechanisms.
This motivates modulating the vanilla attention logits to enlarge the gap(i.e., $\tilde{\gamma}\geq\gamma$), thereby tightening the leakage bounds. We then propose Spatially-Modulated Attention (SMA), a lightweight module that enforces spatial constraints based on mask boundaries to prevent the model from indiscriminately using the entire context to synthesize background, as shown in Fig. 4.
In standard self-attention, logits are $A_{ij}=\frac{Q_{i}K_{j}^{\top}}{\sqrt{d_{k}}}$. To control attention flow across semantically separated regions, SMA adds a mask-conditioned bias using binary region labels $m_{i},m_{j}\!\in\!\{0,1\}$ ($1=$ foreground, $0=$ background):
where $\alpha_{st}$ are four learnable scalars, maintained per self-attention block and shared across its attention heads. The indicator $\mathds{1}[m_{i}=s,m_{j}=t]$ equals $1$ when token $i$ lies in region $s$ and token $j$ in region $t$, and 0 otherwise.
We initialize all biases at zero to preserve vanilla attention at the start and enable adaptive spatial modulation during training. To encourage suppression of cross-region interactions, we enforce hard constraints on the bias terms: $\alpha_{10},\alpha_{01}\!\leq\!0$ (penalizing cross-domain attention) and $\alpha_{11},\alpha_{00}\!\geq\!0$ (reinforcing within-domain attention). These constraints are implemented as hard boundaries during optimization, ensuring consistent suppression of cross-region interactions throughout training.
This design provides explicit, mask-driven control over attention flow without altering the core attention mechanism. Deploying SMA mechanism widens the CAG, cuts leakage exponentially (Theorem 2), and improves training conditioning and stability at minimal cost.
3.5 Optimization Objective
During training, each foreground branch learns to reconstruct its assigned region while the background branch predicts the noise component in the diffusion process. Following standard diffusion model formulation, given a clean image $x_{0}$ and a randomly sampled timestep $t$, we obtain the noisy input $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ where $\epsilon\sim\mathcal{N}(0,I)$ is the target noise. Each branch is trained to predict this target noise within its respective domain:
$\mathcal{L}_{k}=\left\|M_{k}\odot\big(\epsilon-\epsilon_{k}(x_{t},t)\big)\right\|_{2}^{2},\quad\mathcal{L}_{\text{bg}}=\left\|\,\epsilon-\epsilon_{\text{bg}}(x_{t},t)\,\right\|_{2}^{2}.$ (4)
where $\epsilon_{k}(x_{t},t)$ and $\epsilon_{\text{bg}}(x_{t},t)$ denote the noise predictions from the $k$-th foreground branch and background branch respectively, conditioned on the noisy input $x_{t}$ and timestep $t$.
The total loss combines foreground and background terms with a training-step warm-up: